OpenStudy (anonymous):

Find the general solution of: xdy+ydx= (xy)^2 dy

OpenStudy (jamesj):

x dy + y dx = d(xy) Hence \[\frac{ d(xy)}{(xy)^2} = dy\] So you can now integrate both sides

OpenStudy (anonymous):

Okay, so the answer I got is xy=y+c

OpenStudy (jamesj):

Not quite. The integral on the LHS is something else.

OpenStudy (anonymous):

hmm, well if this is not it, then I give up ): is it: -xy=2y+C ?

OpenStudy (jamesj):

What is \[\int\limits \frac{du}{u^2}\]